OpenStudy (anonymous):

A candy store makes a 9-pound mixture of gummy candy, jelly beans and hard candy. The cost of gummy candy is $2.00 per pound, jelly beans cost $3.00 per pound, hard candy cost $3.00 per pound. The mixture calls for two times as many gummy candy piece as jelly beans. The total cost of the mixture is $23.00. How much of each ingredient did the store use?

hero (hero):

g = pounds of gummy j = pounds of jelly h = pounds of hard g + j + h = 9 2g + 3j + 3h = 23 g = 2j

hero (hero):

g + j + h = 9 represents the total pounds of all the candy. 2g + 3j + 3h = 23 represents the total cost of all the candy g = 2j represents that the amount of gummy is twice the amount of jelly.

hero (hero):

Since g = 2j we can substitute 2j in place of g in the first two equations to get: 2j + j + h = 9 2(2j) + 3j + 3j = 23 Combining the j terms together: \(3\color\red{\text{j}} + \color\green{\text{h}} = 9\) \(7\color\red{\text{j}} + 3\color\green{\text{h}} = 23\) Now we have a system of two equations @DestinyMcGee2015 Can you solve the system of two equations?

OpenStudy (anonymous):

Yes, thank you so much !

hero (hero):

Okay, solve it real quick and let me know what you get.